# DiffuseDef: Improved Robustness to Adversarial Attacks

### v1

@@ Appendix A Data Preparation
Table 5 presents the number of examples in train/valid/test splits and the average token length for the three datasets used in the experiments. For QNLI and AGNews datasets, we randomly split the training set into our train/valid splits, with a ratio of 0.9/0.1, and use its test set as our test split. For IMDB dataset, we randomly split the dataset into train/valid/test splits with a ratio of 0.8/0.1/0.1. All train/valid/test splitting is performed using a random seed of 42.
@@ Appendix B Evaluation Constraints
When evaluating with adversarial attack, We follow the parameter settings for TextAttack as suggested in (Li et al., 2021). The minimum semantic similarity between the clean text and adversarial text is set to 0.84, with the score computed using Universal Sentence Encoder (Cer et al., 2018). The maximum number of candidate substitution from attacker is 50, thus the maximum number of queries where is the number of tokens. Finally, the maximum percentage of changed tokens is set to 0.3/0.1/0.2 for AGNews, IMDB, and QNLI dataset respectively.
@@ Appendix C Training
The details on hyper-parameters of diffusion training can be found in Table 7. All models are trained on a single RTX A6000 GPU. The diffusion training of 100 epochs takes 6/4/3 hours on AGNews, IMDB, QNLI datasets respectively.
@@ Appendix D License for Scientific Artifacts
Table 8 lists the scientific artifacts including data, codes, and models used in this paper. The use of these artifacts in this paper is consistent with their intended use, i.e. for scientific research only. The data used in the experiment is in English and does not contain personally identifying info or offensive content.
Appendix E Example of noising and denoising in DiffuseDef
@@
Adding and removing noise to hidden states are essential features in DiffuseDef which contribute to the improved adversarial robustness. To study how adding or removing noise can affect the semantic meaning of the text, we feed the hidden states to the pretrained BERT model with masked language modeling (MLM) head to generate the text output.
@@
In Table 9, we present the MLM outputs from hidden states added with different steps of noise and the MLM outputs from noise hidden states denoised with same number of steps. In the example shown, with more noise added some semantic information can be lost and replaced by symbols or function words like "." or "the". In contrast, denoising for the same number of steps help alleviate such information lost. For example, the word "IBM" can be recovered from the noise.
@@
However, in practise it is not possible to assume number of denoising steps therefore in Table 10 we show the MLM outputs of denoised hidden states directly from clean and adversarial texts. On clean text, we observe that a higher number of denoising steps can result in more abstraction of the texts. For example, more words are replaced with "the" in the MLM outputs as grows. However, words related to the topic (e.g. "Manchester United", "Liverpool") are kept during the denoising process, thus the model can predict correctly. Similarly, the trend of abstraction can be also found on adversarial text while we observe that the denoising can help remove the adversarial noise / perturbation and recover the word "united" from "nation", thus resulting its correct prediction on the adversarial text.
@@
Appendix F Confusion Matrix under Attack
Figure 5 and 6 present the confusion matrixes of models prediction on clean text and on adversarial texts (successful attack example) on AGNews and IMDB test sets respectively.

### v2

@@ Appendix A Data Preparation
Table 6 presents the number of examples in train/valid/test splits and the average token length for the three datasets used in the experiments. For QNLI and AGNews datasets, we randomly split the training set into our train/valid splits, with a ratio of 0.9/0.1, and use their test sets as our test split. For IMDB dataset, we randomly split the dataset into train/valid/test splits with a ratio of 0.8/0.1/0.1. All train/valid/test splitting is performed using a random seed of 42.
@@ Appendix B Evaluation Constraints
When evaluating with adversarial attack, we follow the parameter settings for TextAttack as suggested in (Li et al., 2021). The minimum semantic similarity between the clean text and adversarial text is set to 0.84, with the score computed using Universal Sentence Encoder (Cer et al., 2018). The maximum number of candidate substitutions from the attacker is 50, thus the maximum number of queries where is the number of tokens. Finally, the maximum percentage of changed tokens is set to 0.3/0.1/0.2 for AGNews, IMDB, and QNLI datasets respectively.
@@ Appendix C Training
The details on hyperparameters of diffusion training can be found in Table 8. All models are trained on a single RTX A6000 GPU. The diffusion training of 100 epochs takes 6/4/3 hours on AGNews, IMDB, and QNLI datasets respectively.
@@ Appendix D License for Scientific Artifacts
Table 9 lists the scientific artifacts, including data, codes, and models used in this paper. The use of these artifacts in this paper is consistent with their intended use, i.e., for scientific research only. The data used in the experiment is in English and does not contain personally identifying information or offensive content.
Appendix E Robustness with RoBERTa
In Table 10, we report the black-box attack results on RoBERTa-based models. A similar conclusion holds for BERT-based models, and DiffuseDef significantly outperforms other defense methods on all 3 datasets.
| Dataset | PLM | Method | Clean% | AUA% | #Query | ||||
| TF | TB | BA | TF | TB | BA | ||||
| AGNews | RoBERTa-base | Fine-Tuned | 94.9 | 34.1 | 36.9 | 43.6 | 372 | 396 | 410 |
| InfoBERT | 95.5 | 40.2 | 45.2 | 48.6 | 392 | 421 | 430 | ||
| FreeLB++ | 95.4 | 57.5 | 62.9 | 55.9 | 444 | 467 | 447 | ||
| RSMI | 93.1 | 64.2 | 66.4 | 67.4 | 774 | 861 | 808 | ||
| DiffuseDef-RSMI (Ours) | 92.9 | 82.9 | 83.5 | 82.2 | 905 | 925 | 1047 | ||
| IMDB | RoBERTa-base | Fine-Tuned | 94.6 | 21.3 | 17.9 | 13.6 | 587 | 671 | 493 |
| InfoBERT | 94.8 | 30.9 | 27.9 | 21.8 | 681 | 760 | 549 | ||
| FreeLB++ | 95.3 | 46.0 | 42.1 | 33.9 | 829 | 974 | 637 | ||
| RSMI | 92.7 | 77.9 | 74.3 | 70.6 | 3443 | 4342 | 2619 | ||
| DiffuseDef-RSMI (Ours) | 92.4 | 84.7 | 84.1 | 84.3 | 3673 | 4782 | 3007 | ||
| QNLI | RoBERTa-base | Fine-Tuned | 92.8 | 26.6 | 22.5 | 20.3 | 204 | 219 | 188 |
| InfoBERT | 92.5 | 29.1 | 25.8 | 20.3 | 205 | 223 | 189 | ||
| FreeLB++ | 92.8 | 33.9 | 27.7 | 20 | 227 | 244 | 200 | ||
| RSMI | 89.3 | 38.8 | 33.2 | 30.0 | 340 | 375 | 343 | ||
| DiffuseDef-FreeLB++ (Ours) | 92.7 | 64.6 | 64.3 | 61.5 | 473 | 579 | 524 | ||
| DiffuseDef-RSMI (Ours) | 88.8 | 57.7 | 55.7 | 53.9 | 469 | 578 | 518 | ||
Appendix F Robustness w.r.t Token Length
Figure 5 provides a comparison of the defense rate for different models by token length on the IMDB dataset. The defense rate is calculated as the percentage of test examples in which TextFooler fails to construct a successful attack. All models except RSMI show a consistent trend that the defense rate declines as the texts lengthen. This trend can be attributed to the nature of adversarial attacks, as longer texts allow for the generation of more adversarial examples. Specifically, adversarial training defense methods like InfoBERT and FreeLB++ show poor performance on longer texts (more than 300 tokens), with the defense rate reduced to nearly 0. This drastic decline indicates that, given an adequate number of queries, the attacker is guaranteed to find a successful attack to fool these models. Similarly, EarlyRobust exhibits a performance drop on long texts as it is based on FreeLB training. RSMI, however, performs worse on short texts, but its defense rate increases as the text length grows. Compared to all SOTA defense approaches, the two DiffuseDef variants show a more steadily declining trend and maintain a higher defense rate across all token lengths, i.e., DiffuseDef is more robust to input text length.
Appendix G Example of noising and denoising in DiffuseDef
@@
Adding and removing noise to hidden states are essential features in DiffuseDef, which contribute to the improved adversarial robustness. To study how adding or removing noise can affect the semantic meaning of the text, we feed the hidden states to the pretrained BERT model with a masked language modeling (MLM) head to generate the text output.
@@
In Table 11, we present the MLM outputs from hidden states added with different steps of noise and the MLM outputs from noise hidden states denoised with the same number of steps. In the example shown, with more noise added some semantic information can be lost and replaced by symbols or function words like “.” or “the”. In contrast, denoising for the same number of steps helps alleviate such information loss. For example, the word “IBM” can be recovered from the noise.
@@
However, in practice, it is not possible to assume the number of denoising steps; therefore in Table 12 we show the MLM outputs of denoised hidden states directly from clean and adversarial texts. On clean text, we observe that a higher number of denoising steps can result in more abstraction of the texts. For example, more words are replaced with “the” in the MLM outputs as grows. However, words related to the topic (e.g., “Manchester United”, “Liverpool”) are kept during the denoising process, thus the model can predict correctly. Similarly, the trend of abstraction can also be found on adversarial text while we observe that the denoising can help remove the adversarial noise/perturbation and recover the word “united” from “nation”, thus resulting in its correct prediction on the adversarial text.
@@
Appendix H Confusion Matrix under Attack
Figure 6 and 7 present the confusion matrices of models’ predictions on clean text and on adversarial texts (successful attack example) on AGNews and IMDB test sets respectively.